OpenStudy (anonymous):

derivative: 1. f(t)= (t/t-3)^2 2. h(t) = (t^2/t^3+2)^2

OpenStudy (aum):

f(t)= (t/t-3)^2 To find the derivative, apply the power rule, chain rule and quotient rule.

OpenStudy (anonymous):

f(t) = (t)^2/(t-3)^2 ???then?

OpenStudy (aum):

Use the quotient rule.

OpenStudy (aum):

But I'd use the power rule first and then the quotient rule. \[ f(t) = \left(\frac{t}{t-3}\right)^2 \\ f'(t) = 2\left(\frac{t}{t-3}\right)^1\frac{d}{dt} \left(\frac{t}{t-3}\right) \] Now apply the quotient rule.

OpenStudy (anonymous):

the derivative of (t/t-3) is 0 right?

OpenStudy (aum):

\[ \frac{d}{dt}\left(\frac{u}{v}\right) = \frac{vu'-uv'}{v^2} \\ \frac{d}{dt}\frac{t}{t-3} = \frac{(t-3)(1) - t(1)}{(t-3)^2} = \frac{-3}{(t-3)^2} \]

OpenStudy (aum):

Read about quotient rule in your book if it is not clear.

OpenStudy (anonymous):

f'(t)= 2(t/t-3)x -3/(t-3)^2 then?>

OpenStudy (aum):

\[ f'(t) = \frac{2t}{(t-3)} * \frac{-3}{(t-3)^2} = -\frac{-6t}{(t-3)^3} \]

OpenStudy (anonymous):

can you help me with others?

OpenStudy (anonymous):

\[g(t)=\sqrt{\sqrt{t+1}} +1 \]